A horse canters away from its trainer in a straight line, moving 37 m away in 9.1 s . It then turns abruptly and gallops halfway back in 1.8 s .

Calculate its average speed for the entire trip.

Express your answer using two significant figures.

Calculate its average velocity for the entire trip, using "away from the trainer" as the positive direction.

Express your answer using two significant figures.
Answer #1

Total distance = 37 m + halfway back = 37 m + 18.5 m = 55.5 m
total time = 9.1 s + 1.8 s = 10.9 s
savg = 55.5 m/10.9 s = 5.09 m/s

For velocity, we will calculate displacement instead of distance for speed.
total displacement = 37 m - 18.5 m = 18.5 m
total time = 9.1 s + 1.8 s = 10.9 s
vavg = 18.5/10.9 s = 1.68 m/s